When I came in this morning, I noticed that the Mode Cleaner had not been locked for at least the past 8 hours. We moved the MC SUS sliders until the MC SUSPIT and SUSYAW values for each mirror were back to approximately the place they were the last time the MC was nicely locked (~12 hours ago). This got the MC flashing TEM00, so we thought we were doing well.
However, if the servo was enabled, any time the cavity flashed a small-order mode (especially 00), the mirrors would get super kicked. Not good.
We went to investigate, and discovered that the RFPD aux laser was left on again. We turned that off, however that didn't fix the situation.
Manasa suggested checking that the WFS were really, really off. When we looked at the WFS master screen, we noticed that although the WFS servos were off, the MC mirrors' ASC filter banks had non-zero inputs. We checked, and this is not from the MCASS, nor is it from the MC WFS lockins. At this point, I have no idea where these signals are coming from. I have turned off the ASC outputs for all the MC mirrors (which means that we cannot turn on the WFS), and the MC locks fine.
So, we need to know where the ASC signals are coming from. There isn't anything that I can see, from any screen that I can find, that indicates some signals being sent over there. Has anyone done anything lately? I know Koji was working on IPC stuff the other day, but the MC was locking fine over the weekend until yesterday afternoon, so I suspect that's not the culprit.
I have turned off the outputs of the WFS lockins, as part of my turning things off, so if whatever script needs them doesn't enable them, they should be turned back on by hand.